OpenStudy (anonymous):

Find the equation in standard form which passes through (–1, 4) and (0, 9). You must show all work.

OpenStudy (mimi_x3):

m = y^2 - y^1/x^2 -x^1 = (9-4)/(0--1) = 5/1 = 5 y- y^1 = m(x-x^1) y - 4 = 5(x+1) y-4 = 5x + 5 5x - y + 9 = 0